OpenStudy (anonymous):

Determine the solution set of x^2 - 80 = 0. a) {±2} b) {±} c) {±4}

OpenStudy (anonymous):

so add the 80 over and take the square root, right? The square root of 81 is +/- 9, so you're answer should be close to that. Is that all it gives for b?

OpenStudy (anonymous):

no sorry I posted the options wrong. a) {±2 square root of 5} b) {± square root of 5} c) {±4 square root of 5} @caseyrt

OpenStudy (anonymous):

ah ok. So you've taken the square root of 80, right? But you can divide 80 by a perfect square. Try dividing it by 4, 9 and 16 and see what you get. Then you can factor out that perfect square to the outside and leave the rest inside the root.

OpenStudy (anonymous):

\[\sqrt{80} = \sqrt{16*5} = +/- 4 \sqrt{5}\]